OpenStudy (anonymous):

(z^6x^6)^-3 solve with steps please.

OpenStudy (anonymous):

\[\large (z^6x^6)^{-3}\]Is this the question? It can't be solved. Do you mean simplify?

OpenStudy (anonymous):

yes simplify

OpenStudy (anonymous):

yup that's the equation

OpenStudy (anonymous):

\[(x^a)^b = x^{ab}\]\[(x^ay^b)^c=x^{ac}y^{bc}\]

OpenStudy (anonymous):

would the answer be \[z ^{9}x ^{9}\]

OpenStudy (anonymous):

exponents are multiplied and don't forget the negative sign

OpenStudy (anonymous):

so then would it be \[\frac{ 1 }{ z ^{18}x ^{18} }\]

OpenStudy (anonymous):

yup!!

OpenStudy (anonymous):

so let's say if i have a problem like \[(-2z ^{2})(5z ^{3})\]

OpenStudy (anonymous):

i multiply the exponents?

OpenStudy (anonymous):

@ChmE

OpenStudy (anonymous):

oh! ok so all the problems like that multiply the exponents?

OpenStudy (anonymous):

what do you mean by base?

OpenStudy (anonymous):

\[(-2z ^{2})(5z ^{3})=-10z^5\]@marixoxo76